OpenStudy (anonymous):

2/3 of students work part-time. Of 6,900 students, how many work part-time? i got 4600, is that correct?

OpenStudy (lgbasallote):

2/3 x 6900 = 2 x 2300 = 4600 \[\checkmark\]

OpenStudy (anonymous):

How did you set it up?

OpenStudy (lgbasallote):

what do you mean?

OpenStudy (anonymous):

Where did 2300 come from?

OpenStudy (lgbasallote):

6900/3

OpenStudy (lgbasallote):

\[\frac 23 \times 6900 \implies 2 \div 3 \times 6900 \implies 6900 \div 3 \times 2\] it really doesnt make any difference

OpenStudy (lgbasallote):

i just divided first so i have a smaller number to multiply so it will be easier to compute mentally hehe

OpenStudy (lgbasallote):

anyway bottomline is you're right

OpenStudy (anonymous):

Ok thanks I knew the answer already because they were given to us I just had to find out how to get the answers.. So just divide the denominator by 6900?

OpenStudy (lgbasallote):

\[\frac 23 \times 6900 \implies \frac{2 \times 6900}{3}\] you might wanna do it like that

OpenStudy (lgbasallote):

but yeah..you can also do what i did

OpenStudy (lgbasallote):

all roads lead to Rome so they say

OpenStudy (lgbasallote):

ok i need to go now so i hope you got that :D